# Vital Insight: Assisting Experts’ Context-Driven Sensemaking of Multi-modal Personal Tracking Data Using Visualization and Human-In-The-Loop LLM

### v3

@@ 7.3.4. General Usability and Qualitative Feedback
Regarding trust in AI, experts seemed to trust the AI-generated content more than the first iteration because VI-2 explicitly describes the evidence used and the content aligns with that evidence. While experts built more trust, they also started to “ask for more”, compared to Study 2, where experts were more doubtful about the AI content. For instance, P17 mentioned liking one of the occurrence explanations because it provided an extensive description that combined multiple sensing data points to make a “real” inference, rather than simply extracting user check-in data. This reflects something that wasn’t fully revealed in Study 2: beyond merely verifying AI content, as users build trust in the AI, they treat it as additional modality of information, and even use it to validate other information. In one scenario, P13 observed a user check-in indicating that the person went out for milk tea and the grocery store, but they still looked at the LLM generated content to validate the information. As experts gradually built trust in the LLM and started to view it as another reliable source of information, they spent more time comprehending and integrating its insights.
@@ 7.3.5. User Interaction: Experts’ sensemaking process
In Study 3, we posed specific questions for experts to answer. This approach revealed new insights to the sensemaking process that is complementary to the more open-ended summary-writing scenario in Study 2. Some experts immediately began identifying related data based on the question. The first thing they tried to identify was time, which helped them narrow down the data set to focus on. Some questions also prompted experts to focus on a specific modality. For instance, when asked what the person was doing during a long phone usage period, many experts went directly to the phone usage data and only turned to other modalities as necessary. In such cases, LLM occurrences became an additional modality, on par with other sensing modalities, assisting experts to make sense of what happened. Interestingly, despite the specific questions, some experts still started by reviewing the entire day’s data to first get an overview. This process is similar to what we observed in Study 2. In this scenario, both LLM occurrences and the day-in-a-glance feature were the first things some experts examined, especially when they were familiar with LLM and “knew what to expect.”
@@
Figure 8. Experts employ detailed sensemaking procedures across various modalities, including sensing data, contextual information and AI contents to identify data patterns, construct personas, and determine potential actions.
[…]
Diagram illustrating sensemaking procedures with Sensing Data, Contextual Information, Pattern identification, AI-supported Inference, and Action. Sensing Data like heart rate and steps are compared and integrated into Patterns such as discrepancies and anomalies. Contextual Information, including user profile and location, helps extract or fill and explain patterns. AI-supported Inference covers inference, summary, and questions, interacting with Patterns to guide Actions like stakeholder outreach and debugging. Arrows indicate processes such as compare, integrate, extract or fill, and seek explanation, showing the flow between components. updated figure
@@ 8. Experts’ sensemaking model
Experts frequently compare various sensing modalities and integrate them to synthesize patterns. We identified five types of patterns that experts pay attention to: long duration, change, discrepancy, gap, and routine activities. To explain these patterns, experts turn to contextual information like user profiles and check-ins, and use them to fill in gaps that sensing data alone cannot provide, often involving a process of extracting relevant details. Based on the patterns, experts may take certain actions directly, such as addressing sensor issues. More commonly, however, experts work to synthesize a user persona before taking action. These personas are typically constructed from patterns observed in longitudinal patterns, along with known contextual information such as demographics. Common persona traits include lifestyle, hobbies, social interactions, and regular routines. Once the persona is established, experts use it to guide specific actions, such as reaching out to relevant stakeholders or designing personalized interventions. Furthermore, the sensemaking procedures described above can be supported by AI inference at varying levels of granularity. For instance, observation summaries of sensing data can help experts compare and integrate raw data across modalities to identify patterns, while more creative inferences can assist them in brainstorming potential explanations for those patterns.
@@
To further elevate the sensemaking model, we consider the collaboration between experts, data, and AI. Experts iteratively move between direct data representation and indirect inference to generate insights (Fig. 9). A common approach for experts is to explore the data directly and develop their thoughts. Two other main paths involve indirect inference. One starts with questioning certain aspects of an indirect inference and seeking evidence from data to validate the results. The other involves being inspired by AI-supported inference, such as overlooked details in indirect inference, and then intentionally retrieve relevant information from the data. These three paths lead to insights, where experts feel confident in drawing conclusions. After reaching insights, some experts may revisit the indirect inference to confirm their assumptions. In addition to generating insights, these paths may also lead to confusion. Some experts choose to consult indirect inference for hints while others routed to stakeholders like participants to reach confident answers.
@@
Experts’ sensemaking model also highlights the need for customization. While there are many commonalities in their mental models and thought processes, each expert’s focus varies slightly. Differences include starting points, preferred modalities, the amount of data or time range they are comfortable reviewing, familiarity with the target population, and preferred processing modes (e.g., more visual or more text-based). Based on this model, the system should allow extensive customization across these aspects, including basic interface customization like layout and color following Schneiderman’s mantra “Overview first, zoom and filter, then details-on- demand” (shneiderman2003eyes, ), as well as advanced customization based on the personal sensemaking model. It’s also important to recognize the varying levels of trust in AI-generated content among different experts. Systems should tailor the tone of the AI to be more conservative, with additional proof and explanations for certain experts, while provide more insights from the LLM for those open to getting inspiration from the outputs.
@@
Figure 9. High-level experts sensemaking procedure around direct data representation and AI-supported inference.
[…]
Diagram of the high-level sensemaking procedure with Direct Data Representation, Expert Inference, and AI-supported Inference. Direct Data Representation on the left, expert inference in the middle with two subcomponents insight and confusion, AI supported inference on the right, with an expert icon under expert inference and stakeholder under AI-supported inference. Different components are connected through arrows. AI-supported inference ‘inspires’ experts to ‘retrieve’ data from direct data representation, then ‘generate’ insights as expert inference. Experts ‘question’ AI-supported inference, ‘validate’ from direct data representation, and ‘generate’ insights as expert inference. Experts ‘explore’ direct data representation, and ‘generate’ insights as expert inference. There are also two arrows between expert inference and AI-supported inference, where experts ‘confirm’ their insights from AI, and ‘consult’ AI when there’s confusion. AI-supported inference also points to stakeholders icon.
@@ 9.1.1. Evidence-Based Sensemaking that Enables Validation to Build Trust
The design of these tools must support evidence-based sensemaking, enabling a continuously validating and confirming process by skeptical experts. It should explicitly provide and identify the evidence used to generate indirect inferences, and allow experts to validate these findings using their own sensemaking methods. This aligned with the challenges of self-calibration as outlined by Karahanoğlu et al. (Karahanoğlu_View, ). We further categorized 2 useful evidences in this scenario: source of data being used to conclude the inference, and the explanation of the deduction procedure using these sources. These two evidences highlights experts’ needs of the AI assistance to have a human-like sensemaking process instead of machine-like strategic such as black-box in Machine learning. Moreover, the system should ensure a clear distinction between raw data and inferences to provide clear expectations to facilitate trust.
@@ 9.1.2. Integration of Modalities to Assist Contextualization
Tools must facilitate the seamless integration of various sensing data modalities and enable easy comparison of different modalities across the same scale. This capability enables experts to compare and analyze different data types simultaneously, offering a comprehensive view of an individual’s behaviors. In our cases, time is the most important scale in making sense of multi-modal health sensing data. As a result, visually aligning multiple raw sensing data modalities on a time scale instead of only showing fragmented summary of a single modality is a helpful design decision to facilitate integration. We also noticed that other than raw sensing data that naturally aligned through time, contextual information such as self-checkin conversation and user profile are usually hard to aligned the rest of the data. This is also proved to be a main challenge in our user testing sessions. As a result, an effective tool must provide and extract important contextual insights to help explain events. This includes integrating information about the individual, such as their user profile, proactive self-report data, and calculated features like location labels and baselines. After extracting the information, aligning it back to the same scale of time is very helpful for experts’ sensemaking process. All these efforts help the contextualizing of complex data to generate meaningful insights.
@@ 9.1.3. Support an Iterative Sensemaking Process and Determining Actions
9.1.4. Provide Both AI Assisted Contextual Overview and Momentary Highlights
@@ 9.3. Biases in LLM-generated inferences
During user testing with experts, we observed them gradually building trust in the LLM-generated inferences and summaries while exploring the various system components. However, this process also increased the potential risk of bias introduced by inaccurate LLM results. Several prior works have discussed concerns with bias in LLM-generated outcomes: highlighting biases related to gender (wan2023kelly, ), age (duan2024large, ), geography (manvi2024large, ), and politics (urman2023silence, ). These biases could influence the interpretation of contextual information about individuals. For example, LLM may offer a stereotyped interpretation of the data based on the user’s certain demographics. LLMs also exhibit bias in their information retrieval processes, such as giving preference to documents or items from specific input positions like the beginning or end of the list (dai2024neural, ; li2023halueval, ; dai2024bias, ). This bias could potentially affecting the retrieval of various raw sensing data. Since our user testing primarily involved people with extensive expertise, they expressed a highly skeptical view of the LLM-generated results, often validating the inferences before accepting them. For example, P2 specifically mentioned concerns about being influenced by the LLM summary and stated they would avoid checking it first to prevent bias. However, as their trust in the system grew, there is a risk that they might be increasingly influenced by the LLM’s outputs, even for experts (choudhury2024large, ; kerasidou2022before, ; ryan2020ai, ). During our internal evaluation of the LLM-generated results, we also noted a potential concern with over-interpreting sensor data, where the LLM might notice trivial fluctuations in heart rate and force itself to make an inference. Without explicit instructions, LLMs also tend to not question the validity and reliability of the data and always offer a potential explanation regardless. While experts may be able to identify incorrect outputs from the LLM, there could still be a subtle bias introduced. Even without considering incorrect outputs, the possibilities provided by the LLM can introduce a small bias in what experts perceive as ’what might have happened.’ Future research must recognize the importance of examining both the LLM’s biases in interpreting personal tracking data and the biases experts might develop when integrating those results.
@@ 9.4. Designing LLM agents that mimic experts’ sensemaking process
Although this study did not aim to enhance the accuracy of complex activity recognition and sensemaking, our insights into experts’ sensemaking processes may help inform the design of more effective AI-assisted systems for interpreting personal tracking data. Guiding LLMs with expert knowledge is a common approach for solving domain-specific problems, including techniques like adding instructions in prompts (ji2024hargpt, ; cosentino2024towards, ), creating expert agents (yang2024drhouse, ), and fine-tuning (kim2024health, ; cosentino2024towards, ). In the context of making sense of sensor data, HARGPT employs simple role-playing prompts without expert guidance to identify activities from raw IMU data (ji2024hargpt, ). Health-LLM evaluated zero/few-shot learning and fine-tuning with simple prompts on sensor data (kim2024health, ). Similarly, prior works have also explored how to encode sensor data to prompts: LLMSense and PhysioLLM use text-formatted sensor data and prompts to derive high-level inferences from sensor traces (ouyang2024llmsense, ; fang2024physiollm, ); Cosetino et al. incorporated domain knowledge into prompts and fine-tuned models using expert responses (cosentino2024towards, ). We build on previous works to provide an iterative and evolving approach to continuously include domain expertise as Data Interpretation Guidance in the prompt. While prior works have explored incorporating domain/expert knowledge, we observed that much of the expert knowledge is usually “static.”. From related work, it is clear that the complex sensemaking processe experts use when interpreting personal tracking data are under-documented and rarely considered to guide LLMs. Our study reveals that these sensemaking processes extend beyond basic knowledge, like the normal range of a person’s heart rate, and encompass the selection and integration of multiple data streams, handling data discrepancies and uncertainties, and strategies for information retrieval. These domain-specific insights can guide LLMs in various ways. For instance, understanding how experts initiate sensemaking – starting from changes rather than following a strict chronological order – could be used to optimize data retrieval for LLMs. Additionally, the differing definitions of anomalies between the LLM and experts highlight an additional area for improvement, involving extending the system to incorporate more task-specific and sophisticated anomaly detection algorithms. Our findings on how experts collaborate with LLMs and other informational modules to reach conclusions underscore how LLM detection models should be designed for real-world applications to support experts effectively. We hope that, although not the main focus of this study, our insights will help AI and LLM researchers recognize opportunities to enhance current detection systems and better support complex sensemaking tasks.
@@ 9.5. Replacing or assisting: Experts and AI collaboration in contextual reasoning of sensing data
Previous works comparing LLM results with human input often focus on tasks with well-defined outcomes, such as activity recognition, stress scores, or diagnoses (yang2024drhouse, ; ji2024hargpt, ; kim2024health, ; cosentino2024towards, ; ouyang2024llmsense, ; fang2024physiollm, ). These tasks allow researchers to define clear metrics and conduct quantitative evaluations to claim success of certain techniques proposed in those studies, like prompt engineering techniques, multi-agent pipeline, fine-tuned model, and more. However, for more subjective tasks like summarizing a person’s day, there is no single “right” or “wrong” answer, thus making objective, quantitative comparisons difficult. For example, we used standardized NLP methods to measure the accuracy and consistency of the summaries, and found that expert-written summaries were significantly more diverse than those generated by LLM. However, these metrics are not enough to claim that LLM behaves better than experts. In real-world studies, both types of results are necessary for different applications. Comparative studies provide valuable insights for building fundamental models that support applications requiring automated outputs. For instance, activity recognition results generated by LLMs can be directly used as context for delivering personalized health interventions. Subjective results, on the other hand, are crucial for human-centered applications that involve multiple stakeholders, such as monitoring dashboards for researchers, daily summaries for self-reflection, or even clinical decision making. Adler et al., in their recent paper, proposed the concept of actionable sensing as an alternative to detection research, which brings passive sensing data on behavior and physiology into clinical encounters to improve patient care (adler2024beyond, ). We embrace a similar concept, where LLM’s extensive narrative capabilities create new opportunities for generating high-level content that is valuable for various stakeholders, extending beyond detection.

### v4

@@ 7.3.4. General Usability and Qualitative Feedback
Regarding trust in AI, experts seemed to trust the AI-generated content more than the first iteration because VI-2 explicitly describes the evidence used and the content aligns with that evidence. While experts built more trust, they also started to “ask for more”, compared to Study 2, where experts were more doubtful about the AI content. For instance, P17 mentioned liking one of the occurrence explanations because it provided an extensive description that combined multiple sensing data points to make a “real” inference, rather than simply extracting user check-in data. This reflects something that wasn’t fully revealed in Study 2: beyond merely verifying AI content, as users build trust in the AI, they treat it as an additional modality of information, and even use it to validate other information. In one scenario, P13 observed a user check-in indicating that the person went out for milk tea and to the grocery store, but they still looked at the LLM-generated content to validate the information. As experts gradually built trust in the LLM and started to view it as another reliable source of information, they spent more time comprehending and integrating its insights.
@@ 7.3.5. User Interaction: Experts’ Sensemaking Process
In Study 3, we posed specific questions for experts to answer. This approach revealed new insights into the sensemaking process that are complementary to the more open-ended summary-writing scenario in Study 2. Some experts immediately began identifying related data based on the question. The first thing they tried to identify was time, which helped them narrow down the data set to focus on. Some questions also prompted experts to focus on a specific modality. For instance, when asked what the person was doing during a long phone usage period, many experts went directly to the phone usage data and only turned to other modalities as necessary. In such cases, LLM occurrences became an additional modality, on par with other sensing modalities, assisting experts to make sense of what happened. Interestingly, despite the specific questions, some experts still started by reviewing the entire day’s data to first get an overview. This process is similar to what we observed in Study 2. In this scenario, both LLM occurrences and the day-in-a-glance feature were the first things some experts examined, especially when they were familiar with LLM and “knew what to expect.”
@@
Figure 8. Experts employ detailed sensemaking procedures across various modalities, including sensing data, contextual information and AI contents to identify data patterns, construct personas, and determine potential actions.:
@@ 8. EXPERTS’ SENSEMAKING MODEL
Experts frequently compare various sensing modalities and integrate them to synthesize patterns. We identified five types of patterns that experts pay attention to: long duration, change, discrepancy, gap, and routine activities. To explain these patterns, experts turn to contextual information like user profiles and check-ins, and use them to fill in gaps that sensing data alone cannot provide, often involving a process of extracting relevant details. Based on the patterns, experts may take certain actions directly, such as addressing sensor issues. More commonly, however, experts work to synthesize a user persona before taking action. These personas are typically constructed from patterns observed in longitudinal data, along with known contextual information such as demographics. Common persona traits include lifestyle, hobbies, social interactions, and regular routines. Once the persona is established, experts use it to guide specific actions, such as reaching out to relevant stakeholders or designing personalized interventions. Furthermore, the sensemaking procedures described above can be supported by AI inference at varying levels of granularity. For instance, observation summaries of sensing data can help experts compare and integrate raw data across modalities to identify patterns, while more creative inferences can assist them in brainstorming potential explanations for those patterns.
@@
To further elevate the sensemaking model, we consider the collaboration between experts, data, and AI. Experts iteratively move between direct data representation and indirect inference to generate insights (Fig. 9). A common approach for experts is to explore the data directly and develop their thoughts. Two other main paths involve indirect inference. One starts with questioning certain aspects of an indirect inference and seeking evidence from data to validate the results. The other involves being inspired by AI-supported inference, such as overlooked details in indirect inference, and then intentionally retrieve relevant information from the data. These three paths lead to insights, where experts feel confident in drawing conclusions. After reaching insights, some experts may revisit the indirect inference to confirm their assumptions. In addition to generating insights, these paths may also lead to confusion. Some experts choose to consult indirect inference for hints while others route to stakeholders like participants to reach confident answers.
@@
Experts’ sensemaking model also highlights the need for customization. While there are many commonalities in their mental models and thought processes, each expert’s focus varies slightly. Differences include starting points, preferred modalities, the amount of data or time range they are comfortable reviewing, familiarity with the target population, and preferred processing modes (e.g., more visual or more text-based). Based on this model, the system should allow extensive customization across these aspects, including basic interface customization like layout and color following Schneiderman’s mantra “Overview first, zoom and filter, then details-on- demand” (shneiderman2003eyes, ), as well as advanced customization based on the personal sensemaking model. It’s also important to recognize the varying levels of trust in AI-generated content among different experts. Systems should tailor the tone of the AI to be more conservative, with additional proof and explanations for certain experts, while providing more insights from the LLM for those open to getting inspiration from the outputs.
@@
Figure 9. High-level experts’ sensemaking procedure around direct data representation and AI-supported inference.:
@@ 9.1.1. Evidence-Based Sensemaking that Enables Validation to Build Trust
The design of these tools must support evidence-based sensemaking, enabling a continuously validating and confirming process by skeptical experts. It should explicitly provide and identify the evidence used to generate indirect inferences, and allow experts to validate these findings using their own sensemaking methods. This aligned with the challenges of self-calibration as outlined by Karahanoğlu et al. (Karahanoğlu_View, ). We further categorized 2 useful pieces of evidence in this scenario: the source of data being used to conclude the inference, and the explanation of the deduction procedure using these sources. These two evidences highlights experts’ need for AI assistance to have a human-like sensemaking process instead of machine-like strategies such as black-box in Machine learning. Moreover, the system should ensure a clear distinction between raw data and inferences to provide clear expectations to facilitate trust.
@@ 9.1.2. Integration of Modalities to Assist Contextualization
Tools must facilitate the seamless integration of various sensing data modalities and enable easy comparison of different modalities across the same scale. This capability enables experts to compare and analyze different data types simultaneously, offering a comprehensive view of an individual’s behaviors. In our cases, time is the most important scale in making sense of multi-modal health sensing data. As a result, visually aligning multiple raw sensing data modalities on a time scale instead of only showing a fragmented summary of a single modality is a helpful design decision to facilitate integration. We also noticed that other than raw sensing data that naturally aligns through time, contextual information such as self-checkin conversation and user profile are usually hard to align with the rest of the data. This is also proven to be a main challenge in our user testing sessions. As a result, an effective tool must provide and extract important contextual insights to help explain events. This includes integrating information about the individual, such as their user profile, proactive self-report data, and calculated features like location labels and baselines. After extracting the information, aligning it back to the same scale of time is very helpful for experts’ sensemaking process. All these efforts help contextualize complex data to generate meaningful insights.
@@ 9.1.3. Support an Iterative Sensemaking Process and Determining Actions
9.1.4. Provide Both AI-Assisted Contextual Overview and Momentary Highlights
@@ 9.3. Biases in LLM-Generated Inferences
During user testing with experts, we observed them gradually building trust in the LLM-generated inferences and summaries while exploring the various system components. However, this process also increased the potential risk of bias introduced by inaccurate LLM results. Several prior works have discussed concerns with bias in LLM-generated outcomes: highlighting biases related to gender (wan2023kelly, ), age (duan2024large, ), geography (manvi2024large, ), and politics (urman2023silence, ). These biases could influence the interpretation of contextual information about individuals. For example, LLM may offer a stereotyped interpretation of the data based on the user’s certain demographics. LLMs also exhibit bias in their information retrieval processes, such as giving preference to documents or items from specific input positions like the beginning or end of the list (dai2024neural, ; li2023halueval, ; dai2024bias, ). This bias could potentially affect the retrieval of various raw sensing data. Since our user testing primarily involved people with extensive expertise, they expressed a highly skeptical view of the LLM-generated results, often validating the inferences before accepting them. For example, P2 specifically mentioned concerns about being influenced by the LLM summary and stated they would avoid checking it first to prevent bias. However, as their trust in the system grew, there is a risk that they might be increasingly influenced by the LLM’s outputs, even for experts (choudhury2024large, ; kerasidou2022before, ; ryan2020ai, ). During our internal evaluation of the LLM-generated results, we also noted a potential concern with over-interpreting sensor data, where the LLM might notice trivial fluctuations in heart rate and force itself to make an inference. Without explicit instructions, LLMs also tend not to question the validity and reliability of the data and always offer a potential explanation regardless. While experts may be able to identify incorrect outputs from the LLM, there could still be a subtle bias introduced. Even without considering incorrect outputs, the possibilities provided by the LLM can introduce a small bias in what experts perceive as ’what might have happened.’ Future research must recognize the importance of examining both the LLM’s biases in interpreting personal tracking data and the biases experts might develop when integrating those results.
@@ 9.4. Designing LLM Agents that Mimic Experts’ Sensemaking Process
Although this study did not aim to enhance the accuracy of complex activity recognition and sensemaking, our insights into experts’ sensemaking processes may help inform the design of more effective AI-assisted systems for interpreting personal tracking data. Guiding LLMs with expert knowledge is a common approach for solving domain-specific problems, including techniques like adding instructions in prompts (ji2024hargpt, ; cosentino2024towards, ; choube2025gloss, ), creating expert agents (yang2024drhouse, ), and fine-tuning (kim2024health, ; cosentino2024towards, ). In the context of making sense of sensor data, HARGPT employs simple role-playing prompts without expert guidance to identify activities from raw IMU data (ji2024hargpt, ). Health-LLM evaluated zero/few-shot learning and fine-tuning with simple prompts on sensor data (kim2024health, ). Similarly, prior works have also explored how to encode sensor data to prompts: LLMSense and PhysioLLM use text-formatted sensor data and prompts to derive high-level inferences from sensor traces (ouyang2024llmsense, ; fang2024physiollm, ); Cosetino et al. incorporated domain knowledge into prompts and fine-tuned models using expert responses (cosentino2024towards, ). We build on previous works to provide an iterative and evolving approach to continuously include domain expertise as Data Interpretation Guidance in the prompt. While prior works have explored incorporating domain/expert knowledge, we observed that much of the expert knowledge is usually “static.”. From related work, it is clear that the complex sensemaking process experts use when interpreting personal tracking data is under-documented and rarely considered to guide LLMs. Our study reveals that these sensemaking processes extend beyond basic knowledge, like the normal range of a person’s heart rate, and encompass the selection and integration of multiple data streams, handling data discrepancies and uncertainties, and strategies for information retrieval. These domain-specific insights can guide LLMs in various ways. For instance, understanding how experts initiate sensemaking – starting from changes rather than following a strict chronological order – could be used to optimize data retrieval for LLMs. Additionally, the differing definitions of anomalies between the LLM and experts highlight an additional area for improvement, involving extending the system to incorporate more task-specific and sophisticated anomaly detection algorithms. Our findings on how experts collaborate with LLMs and other informational modules to reach conclusions underscore how LLM detection models should be designed for real-world applications to support experts effectively. We hope that, although not the main focus of this study, our insights will help AI and LLM researchers recognize opportunities to enhance current detection systems and better support complex sensemaking tasks.
@@ 9.5. Replacing or Assisting: Experts and AI Collaboration in Contextual Reasoning of Sensing Data
Previous works comparing LLM results with human input often focus on tasks with well-defined outcomes, such as activity recognition, stress scores, or diagnoses (yang2024drhouse, ; ji2024hargpt, ; kim2024health, ; cosentino2024towards, ; ouyang2024llmsense, ; fang2024physiollm, ). These tasks allow researchers to define clear metrics and conduct quantitative evaluations to claim success of certain techniques proposed in those studies, like prompt engineering techniques, multi-agent pipeline, fine-tuned model, and more. However, for more subjective tasks like summarizing a person’s day, there is no single “right” or “wrong” answer, thus making objective, quantitative comparisons difficult. For example, we used standardized NLP methods to measure the accuracy and consistency of the summaries, and found that expert-written summaries were significantly more diverse than those generated by LLM. However, these metrics are not enough to claim that LLM behaves better than experts. In real-world studies, both types of results are necessary for different applications. Comparative studies provide valuable insights for building fundamental models that support applications requiring automated outputs. For instance, activity recognition results generated by LLMs can be directly used as context for delivering personalized health interventions. Subjective results, on the other hand, are crucial for human-centered applications that involve multiple stakeholders, such as monitoring dashboards for researchers, daily summaries for self-reflection, or even clinical decision making. Adler et al., in their recent paper, proposed the concept of actionable sensing as an alternative to detection research, which brings passive sensing data on behavior and physiology into clinical encounters to improve patient care (adler2024beyond, ). We embrace a similar concept, where LLMs’ extensive narrative capabilities create new opportunities for generating high-level content that is valuable for various stakeholders, extending beyond detection.